# Personalized Item Representations in Federated Multimodal Recommendation

### v2

@@ 4.6. Complexity Analysis
4.7. Privacy Preservation
FedMR uses a distributed architecture based on FL to effectively reduce the risk of user privacy leakage. Except for the Mixing Feature Fusion Module, all other components of FedMR have the same privacy risk level as existing ID-based federated recommendation methods. In the Mixing Feature Fusion Module, only the locally updated ID embeddings $\mathbf{D}^{(u)}$ and the local accumulated gradients ${\nabla^{(u)}_{j}\gamma}$ are sent to the server, without revealing any raw user data, also maintaining the same communication privacy protection strategy as current methods (Li et al., 2023a; Zhang et al., 2023c; Li et al., 2024b; Feng et al., 2024; Wu et al., 2024). To protect user privacy, randomly selecting participants in each communication round and avoiding the same participants in consecutive rounds can effectively reduce the risk of privacy leakage (Chai et al., 2020; Li et al., 2023a). Additionally, FedMR can incorporate techniques such as differential privacy (Abadi et al., 2016) and stochastic gradient clipping to further enhance privacy protection. We also conduct an ablation study to evaluate the effectiveness of FedMR when additional privacy-preserving measures are applied.
5. Experiments
5.1. Datasets
In this study, we conducte comprehensive experiments on four downstream datasets from Ninerec [^1] The dataset publicly available on: https://github.com/westlake-repl/NineRec (Zhang et al., 2024a) to evaluate the performance and effectiveness of the proposed method FedMR: KU, Bili_Food (Food), Bili_Dance (Dance), and Bili_Movie (Movie). Table 1 provides the detailed statistics of these datasets. The KU dataset is derived from KuaiShou, a well-known short video platform in China, where each entry represents a short video. The Food, Dance, and Movie datasets are collected from the BiliBili platform, focusing on food and cooking videos, dance-related videos, and movie-related videos, respectively. Each dataset presents unique content types and user interaction patterns, but all share a common challenge of high sparsity, with sparsity levels exceeding 95%. For our experiments, we include only users who had interacted with at least 5 items in each dataset to avoid the cold-start problem. Additionally, for missing images in the datasets, we fill them with the average values of visual features; for missing titles, we set ”The title is missing.”
@@ 5.1. Datasets
FedAvg w/ FedMR 0.5147 (356.70% $\uparrow$) 0.1491 (165.30% $\uparrow$) 0.1007 (28.77% $\uparrow$) 0.0261 (29.85% $\uparrow$) 0.0948 (36.21% $\uparrow$) 0.0268 (40.31% $\uparrow$) 0.0564 (2.17% $\uparrow$) 0.0160 (1.27% $\uparrow$)
@@
FCF w/ FedMR 0.5539 (145.63% $\uparrow$) 0.1597 (148.37% $\uparrow$) 0.0699 (3.40% $\uparrow$) 0.0182 (2.75% $\uparrow$) 0.0918 (11.95% $\uparrow$) 0.0247 (12.79% $\uparrow$) 0.0667 (14.80% $\uparrow$) 0.0180 (10.44% $\uparrow$)
@@
FedNCF w/ FedMR 0.1961 (66.75% $\uparrow$) 0.0609 (144.58% $\uparrow$) 0.0708 (220.36% $\uparrow$) 0.0187 (240.00% $\uparrow$) 0.0843 (53.83% $\uparrow$) 0.0243 (49.08% $\uparrow$) 0.0406 (119.46% $\uparrow$) 0.0125 (131.48% $\uparrow$)
@@
PFedRec w/ FedMR 0.2157 (238.62% $\uparrow$) 0.0616 (167.83% $\uparrow$) 0.0392 (55.56% $\uparrow$) 0.0092 (35.29% $\uparrow$) 0.0836 (155.66% $\uparrow$) 0.0207 (122.58% $\uparrow$) 0.0413 (142.94% $\uparrow$) 0.0132 (214.29% $\uparrow$)
@@
Fed FedRAP w/ FedMR 0.4824 (2.51% $\uparrow$) 0.1305 (20.17% $\uparrow$) 0.0895 (31.19% $\uparrow$) 0.0247 (32.91% $\uparrow$) 0.1142 (64.08% $\uparrow$) 0.0294 (53.93% $\uparrow$) 0.0586 (29.36% $\uparrow$) 0.0199 (46.32% $\uparrow$)
5.2. Baselines
To comprehensively evaluate the performance of FedMR, we select 3 cutting-edge centralized MMRec methods, and 5 ID-based federated methods commonly used in FedRec tasks with implicit feedback to serve as baselines in our study:
VBPR (He and McAuley, 2016) incorporates visual features into personalized ranking to improve recommendation accuracy based on implicit feedback.
BM3 (Zhou et al., 2023a) effectively integrates multimodal features through a self-supervised learning framework.
MGCN (Yu et al., 2023) utilizes a multi-view graph convolutional network to separately capture the multimodal features of users and items.
FedAvg (McMahan et al., 2017) is the earliest federated learning method, performing weighted aggregation based on the relative sizes of local client data.
FCF (Ammad-Ud-Din et al., 2019) introduces collaborative filtering into federated learning, assigning equal weights to all clients for average aggregation.
FedNCF (Perifanis and Efraimidis, 2022) updates user embeddings locally on each client and synchronizes item embeddings on the server for global updates.
PFedRec (Zhang et al., 2023c) aggregates item embeddings from clients on the server side and uses them during evaluation to achieve bilateral personalized recommendations.
FedRAP (Li et al., 2023a) achieves fine-grained personalized modeling of users and items through additive personalization.
The methods VBPR, BM3, and MGCN are three centralized benchmark MMRec methods that demonstrate baseline performances of the used datasets in a centralized recommendation setting. The methods FedAvg, FCF, FedNCF, PFedRec, and FedRAP are five ID-based FedRec methods. We conducted experiments on both the vanilla versions of these ID-based FedRec methods and their multimodal versions integrated with our proposed method FedMR to demonstrate the effectiveness of FedMR. Implementations of all used baseline methods are publicly available through their papers. All methods are implemented using PyTorch (Paszke et al., 2019), and all experiments are run on a machine with a 3.60 GHz Intel Xeon Gold 6244 CPU, 3 Quadro RTX 6000 GPUs, and 395 GB of RAM.
5.3. Experimental Setting
Following previous studies (Zhang et al., 2023c; Li et al., 2023a), we perform negative sampling for each positive sample in the training set and use the leave-one-out strategy for validation. We conduct hyperparameter tuning for all baseline methods on each dataset used, selecting the learning rate $\eta$ and adjusting the hyperparameters in FedRAP’s objective function from $\{10^{i}\mid i=-4,\ldots,-1\}$. To ensure fairness, we use the CLIP (Radford et al., 2021a) implemented by the Open CLIP (Ilharco et al., 2021; Radford et al., 2021b; Schuhmann et al., 2022) as the foundation model in our experiments, loading the ViT-B-32 parameters trained by OpenAI to process multimodal information. It should be noted that the foundation model in FedMR can be replaced with other models as needed. Additionally, we add a feature mapping layer before the input fusion layer of the multimodal embeddings and ID embeddings, mapping the feature dimensions output by the foundation model to a fixed latent embedding dimension, i.e. 64, to improve storage efficiency. We fix the training batch size at 2048 for all methods, and for methods with hidden layers, the number of layers is set to 3. For each federated baseline, we adopt the same aggregation strategy to update the global part of FedMR. In the main experiments, all clients participated in the training of both the original versions and the FedMR versions of all federated baselines, and no additional privacy protection measures are applied. Additionally, for these federated baselines, the number of local training epochs was set to 10. Except for loading the foundation model, we do not use any pre-training strategies for any methods.
5.4. Evaluation
In our experiments, we treat the observed interactions as positive samples and the rest as negative samples. We sort the predicted interaction scores between users and all candidate items in descending order, masking the observed positive samples from the training set. Then, we evaluate the recommendation performance using two widely-used metrics: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020). In the experiments, we set $K=50$.
@@ 5.5. Comparison Analysis
Table 2 shows the performance of both centralized and federated methods across the four datasets. The results highlight that FedMR significantly boosts the performance of the used FedRec methods, particularly in the KU dataset, where HR@50 for FedAvg, FCF, PFedRec, and FedRAP exhibits notable improvements. This indicates that FedMR effectively enhances traditional ID-based FedRec methods by integrating multimodal item information. The performance improvements in other datasets further demonstrate FedMR’s ability to alleviate data sparsity by better leveraging users’ interaction histories alongside multimodal data for more personalized recommendations, while protecting user privacy. Compared to centralized MMRec methods, FedRec approaches using FedMR achieve similar or even superior performance while maintaining privacy, e.g., in the KU dataset, FedAvg with FedMR surpasses all centralized methods in terms of HR@50. Federated methods with FedMR also perform strongly across other datasets, reinforcing FedMR’s capability to improve recommendation quality without sacrificing user privacy.
Table 2 further shows that by incorporating multimodal information, FedMR provides a more comprehensive representation of item features. Compared to ID-based FedRec, the use of multimodal data enables more accurate predictions of user preferences. Additionally, the MFFM of FedMR dynamically adjusts the weights of visual, textual, and ID features based on each user’s dependence on these modalities. This flexible approach to feature fusion avoids the drawbacks of single fusion strategy methods, allowing the model to better capture complex relationships between modalities and further enhance personalized recommendation accuracy.
@@ 5.6. Ablation Study
To validate the effectiveness of each component in FedMR, we conducted experiments using several variants of the system. Specifically, we independently applied the Sum, MLP, and Gate strategies for the fusion of multimodal features and ID features, to assess the impact of dynamic weighted fusion based on users’ interaction history $\mathbf{r}_{u}$, $\forall u\in\mathcal{U}$. Additionally, we introduce random Gaussian noise into the gradients uploaded to the server to evaluate the potential performance impact of this privacy-preserving measure. We also examine the distribution of multimodal and ID features within the same feature space. These ablation studies demonstrate the effectiveness of each component of FedMR.
5.6.1. Ablation Study on Modality Distributions
To investigate the distribution relationship between multimodal features and ID features, we select FedRAP (Li et al., 2023a) as the backbone for FedMR in this study. FedRAP maintains both globally shared and user-specific ID features, which helps demonstrate how different ID features affect the distribution within the feature space.
Fig. 2 visualizes the distribution of modality features and ID features after training FedRAP with FedMR, on the KU dataset. The first subfigure in Fig. 2 shows that when the multimodal features of items and globally shared ID features are mapped into a shared representation space, they remain in distinct subregions. This phenomenon, known as the ”Modality Gap” (Liang et al., 2022; Zhang et al., 2023a), helps the model distinguish between different modalities, reducing noise interference and improving the robustness of multimodal data processing. The following two subfigures incorporate the personalized ID embeddings of the users 00 and $42$, respectively, highlighting the significant impact of personalized ID features on the distribution in the shared space. By comparing the first subfigure with the subsequent two, we observe that the introduction of personalized ID features causes the ID features to gradually expand towards the regions of visual and textual features, shrinking the modality gap. This facilitates the fusion of multimodal information and enhances the accuracy of personalized recommendations. Furthermore, by comparing the differences in modality gap changes for each user, it becomes evident that the model can adjust the feature distribution based on user-specific behaviors and preferences, thus better satisfying individual user needs.
This study underscores the importance of personalized ID features in federated recommendation systems. Due to privacy constraints, the model cannot access raw user data directly, but through the use of personalized ID features, the model can provide personalized recommendation services while maintaining user privacy.
Table 3. Performance comparison of FedRec Methods on HR@50 and NDCG@50 with and without multimodal and ID features on the KU dataset.
Method Metric FedMR w/o $\mathbf{V}$ w/o $\mathbf{C}$ w/o $\mathbf{D}$
@@ 5.6.1. Ablation Study on Modality Distributions
FCF HR@50 0.5539 0.1569 (71.67% $\downarrow$) 0.1569 (71.67% $\downarrow$) 0.1275 (76.99% $\downarrow$)
NDCG@50 0.1597 0.0486 (69.57% $\downarrow$) 0.0476 (70.19% $\downarrow$) 0.0287 (82.04% $\downarrow$)
@@
PFedRec HR@50 0.2157 0.2010 (6.82% $\downarrow$) 0.1225 (43.21% $\downarrow$) 0.0490 (77.27% $\downarrow$)
NDCG@50 0.0616 0.0557 (9.58% $\downarrow$) 0.0341 (44.64% $\downarrow$) 0.0191 (68.97% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.2647 (30.78% $\downarrow$) 0.1863 (51.29% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0734 (43.75% $\downarrow$) 0.0490 (62.43% $\downarrow$)
@@ 5.6.2. Ablation Study on Modality Efficiency
Table 3 presents a performance comparison of the five used ID-based FedRec methods with the inclusion of multimodal information. It evaluates these methods under the full model (FedMR) and with the removal of visual modality (w/o V), textual modality (w/o C), and ID features (w/o D), focusing on HR@50 and NDCG@50 metrics.
Firstly, removing the visual modality results in a noticeable performance drop across all methods, particularly for FCF (HR@50 decreased by 71.67%) and FedRAP (HR@50 decreased by 30.78%). This highlights the critical role of visual information in RecSys, especially for FCF. Visual features provide rich, intuitive image-based information that helps models better capture user preferences regarding the appearance of items. Without this visual context, models struggle to accurately infer user preferences, leading to a significant decline in recommendation performance. Secondly, removing the ID feature also causes a substantial decline in performance. The ID feature is essential for personalized recommendations, as it captures users’ historical interactions and reflects their past behaviors and interests. When the ID feature is removed, the model loses its primary mechanism for personalization, resulting in decreased accuracy and personalization of recommendations. This impact is especially pronounced in methods like PFedRec (HR@50 decreased by 82.82%) and FedRAP (HR@50 decreased by 51.92%), which heavily rely on ID features. In contrast, removing the text modality has a relatively smaller impact, though it still affects some methods. For instance, in FedAvg and PFedRec, removing the text modality led to HR@50 decreases of 27.63% and 43.21%, respectively. This indicates that text information plays an important role in capturing semantic relationships between user behaviors and item descriptions, helping the model better understand user needs and item characteristics, thereby enhancing recommendation performance.
In summary, all methods experienced performance degradation to varying degrees when any modality was removed. Considering Fig. 2, visual, text, and ID features are often complementary. Removing any one of these disrupts the balance, making it difficult for the model to fully represent user preferences and item characteristics. Additionally, different methods depend on different modalities to varying extents. For instance, FCF and PFedRec are more dependent on visual information, whereas FedAvg is less reliant on ID features. These differences in dependency lead to varying levels of performance decline when specific modalities are removed. Further experimental results are provided in the appendix.
Figure 3. Performance comparison of different feature fusion strategies in various FedRec methods combined with FedMR on the KU dataset.
5.6.3. Ablation Study on Fusion Strategy
Fig. 3 presents the HR@50 and NDCG@50 evaluation results for five used FedRec methods using different feature fusion strategies with FedMR on the KU dataset. The results indicate that different fusion strategies significantly impact performance. Notably, the Mix strategy, which dynamically combines multiple fusion methods based on user interaction history, outperforms others, while the Sum strategy shows poor performance. This trend is consistent across all metrics.
The Sum strategy simply adds different modality features without considering their complex interactions and inter-relationships. Since each modality’s importance varies, simple addition can lead to information loss or imbalanced weighting, making it difficult to capture relationships effectively. Additionally, RecSys often face a long-tail distribution, where a few popular items dominate user interactions (Park and Tuzhilin, 2008; Yin et al., 2012). The simplicity of the Sum strategy may lead to overfitting on popular items, resulting in poor performance on less interacted items and affecting overall model performance.
In contrast, the MLP strategy uses fully connected networks to learn nonlinear modality combinations, which better captures complex interactions between item modalities and ID features. However, its effectiveness is limited by network depth and the number of parameters. On the other hand, the Gate strategy learns adaptive weights to dynamically filter the contribution of each modality, optimizing the importance of different features.
The Mix strategy integrates all fusion methods and adjusts their weights dynamically based on user interaction history, allowing the model to capture more fine-grained personalized item representations. This approach enables more accurate matching of user preferences and effectively captures complex relationships between multimodal features, leading to superior performance.
This study highlights the importance of personalized item embeddings in FedRec. In FedMR, each client dynamically adjusts fusion strategy weights through the Mixing Feature Fusion Module, based on user interaction history, to generate personalized item embeddings, enhancing personalization and accuracy while maintaining user privacy, thus improving overall system performance.

### v3

@@ 5.1. Datasets
5.2. Baselines
To evaluate the performance of FedMR comprehensively, we compare it with three state-of-the-art centralized MMRecs and five widely used ID-based FedRecs for recommendation tasks with implicit feedback: (1) VBPR (He and McAuley, 2016): Incorporates visual features into personalized ranking to enhance recommendation accuracy for implicit feedback. (2) BM3 (Zhou et al., 2023a): Integrates multimodal features using a self-supervised learning framework. (3) MGCN (Yu et al., 2023): Employs a multi-view graph convolutional network to capture multimodal features of users and items. (4) FedAvg (McMahan et al., 2017): The foundational federated learning method, using weighted aggregation based on local client data sizes. (5) FCF (Ammad-Ud-Din et al., 2019): Combines collaborative filtering with federated learning, performing equal-weighted average aggregation across clients. (6) FedNCF (Perifanis and Efraimidis, 2022): Updates user embeddings locally on clients and synchronizes item embeddings globally on the server. (7) PFedRec (Zhang et al., 2023c): Aggregates client-side item embeddings on the server to enable bilateral personalized recommendations. (8) FedRAP (Li et al., 2023a): Achieves fine-grained personalization for users and items through additive modeling.
The centralized MMRec methods VBPR, BM3 and MGCN serve as benchmarks to gauge the performance potential achievable without federated constraints, highlighting the gap federated approaches aim to close. The selected four federated methods represent common ID-based FedRec approaches. We compare the vanilla versions of the ID-based FedRec baselines against the versions enhanced with FedMR (w/ FedMR) to directly measure the performance gains provided by our framework’s multimodal capabilities [^2] We do not include direct comparisons with AMMFRS (Feng et al., 2024) and FedMMR (Li et al., 2024a) as their official code implementations were not publicly available at the time of our experiments, preventing a fair and reproducible comparison..
5.3. Experimental Setting
Following prior work (Zhang et al., 2023c; Li et al., 2023a), we employ negative sampling during training and use the leave-one-out strategy for validation and testing. Hyperparameters for all baselines are tuned via grid search on the validation set, including the learning rate $\eta$ within $\{10^{i}\mid i=-4,\ldots,-1\}$ and method-specific parameters. We use CLIP (Radford et al., 2021a) as the common foundation model for extracting visual and textual features, implemented via OpenCLIP (Ilharco et al., 2021) with pre-trained ViT-B-32 weights from OpenAI (Radford et al., 2021b; Schuhmann et al., 2022). A linear mapping layer projects multimodal and ID embeddings to a shared latent dimension of 64 for efficiency. The training batch size is 2048. All models incorporating hidden layers use a three-layer MLP structure. For federated methods, clients perform $5$ local training epochs per communication round using a consistent aggregation strategy for both vanilla and FedMR-integrated versions.
5.4. Evaluation
In our experiments, we treat the observed interactions as positive samples and the rest as negative samples. We sort the predicted scores between users and all candidate items in descending order, masking the observed positive samples from the training set. Then, we evaluate the recommendation performance using two widely-used metrics for implicit feedbacks: Hit Rate (HR@K) and Normalized Discounted Cumulative Gain (NDCG@K), which are formally defined in (He et al., 2020). In the experiments, we set $K=50$.
@@ 5.4. Evaluation
FedAvg w/ FedMR 0.5147 0.1491 0.1007 0.0261 0.0948 0.0268 0.0564 0.0160
Impr. 356.70% $\uparrow$ 165.30% $\uparrow$ 28.77% $\uparrow$ 29.85% $\uparrow$ 36.21% $\uparrow$ 40.31% $\uparrow$ 2.17% $\uparrow$ 1.27% $\uparrow$
@@
FCF w/ FedMR 0.5539 0.1597 0.0699 0.0182 0.0918 0.0247 0.0667 0.0180
Impr. 145.63% $\uparrow$ 148.37% $\uparrow$ 3.40% $\uparrow$ 2.75% $\uparrow$ 11.95% $\uparrow$ 12.79% $\uparrow$ 14.80% $\uparrow$ 10.44% $\uparrow$
@@
FedNCF w/ FedMR 0.1961 0.0609 0.0708 0.0187 0.0843 0.0243 0.0406 0.0125
Impr. 66.75% $\uparrow$ 144.58% $\uparrow$ 220.36% $\uparrow$ 240.00% $\uparrow$ 53.83% $\uparrow$ 49.08% $\uparrow$ 119.46% $\uparrow$ 131.48% $\uparrow$
@@
PFedRec w/ FedMR 0.2157 0.0616 0.0392 0.0092 0.0836 0.0207 0.0413 0.0132
Impr. 238.62% $\uparrow$ 167.83% $\uparrow$ 55.56% $\uparrow$ 35.29% $\uparrow$ 155.66% $\uparrow$ 122.58% $\uparrow$ 142.94% $\uparrow$ 214.29% $\uparrow$
@@
Fed FedRAP w/ FedMR 0.4824 0.1305 0.0895 0.0247 0.1142 0.0294 0.0586 0.0199
Impr. 2.51% $\uparrow$ 20.17% $\uparrow$ 31.19% $\uparrow$ 32.91% $\uparrow$ 64.08% $\uparrow$ 53.93% $\uparrow$ 29.36% $\uparrow$ 46.32% $\uparrow$
@@ 5.5. Comparison Analysis
The performance comparison detailed in Table 2 reveals key insights into FedMR’s effectiveness relative to both ID-based federated baselines and centralized multimodal systems. The results consistently demonstrate FedMR’s significant value in bridging the gap between privacy-preserving FL and effective multimodal recommendation.
Enhancing ID-based FedRec. FedMR consistently and substantially elevates the performance of all evaluated ID-based FedRecs across the four datasets. The degree of improvement varies depending on the baseline architecture. Performance gains are particularly dramatic for methods initially exhibiting lower performance, such as FedAvg, FedNCF, and PFedRec, especially on datasets where multimodal signals prove highly beneficial like KU and Food. This trend suggests FedMR effectively integrates crucial multimodal context and personalized fusion capabilities that compensate for the limitations of simpler ID-based approaches in content understanding and heterogeneity handling. While FedMR also enhances more sophisticated baselines like FedRAP, the relative improvement tends to be more moderate, indicating diminishing returns when the base model already incorporates strong personalization mechanisms, though the added multimodal dimension still provides value.
Performance Relative to Centralized Methods. When compared against centralized MMRec benchmarks, FedMR-enhanced methods demonstrate encouraging results, significantly narrowing the typical performance gap observed between federated and centralized systems. Notably, on the KU dataset, FedMR integrated with FedAvg and FCF achieves performance surpassing the centralized counterparts. This suggests that FedMR’s adaptive client-side fusion can be highly effective, potentially adapting better to user nuances than global centralized fusion in certain scenarios. However, a performance gap often persists on the larger Dance, Food, and Movie datasets compared to the best centralized methods. This trend indicates that while FedMR offers substantial improvements, the inherent challenges under the federated recommmendation settings, such as data fragmentation and privacy constraints limiting information sharing, still make fully matching centralized performance difficult, particularly on larger-scale, complex datasets. Furthermore, the lower absolute scores achieved by all methods on datasets like Dance and Movie suggest these represent intrinsically more challenging recommendation tasks, possibly due to complex user behavior or less discriminative multimodal features.
Summary. In summary, FedMR’s integration of multimodal information through the server-client architecture and the adaptive MFFM consistently improves upon diverse ID-based FedRecs. FedMR significantly narrows, and sometimes closes or even surpasses, the performance gap relative to centralized MMRecs. The remaining gap on more challenging datasets highlights the ongoing need for advancements in FedMMRec. FedMR represents a crucial step in this direction, showcasing the effectiveness of adaptive, personalized fusion mechanisms in unlocking multimodal data’s potential under privacy-preserving federated constraints. Specifically, its ability to dynamically adapt to user preferences underscores its value for developing robust and personalized FedMMRecs.
@@ 5.6. Ablation Study
To comprehensively evaluate FedMR’s contributions and key design choices, we conduct ablation studies analyzing four crucial aspects. We investigate the effectiveness of the MFFM compared to simpler alternatives, assess the individual contributions of visual, textual, and ID modalities, examine the impact of client heterogeneity related to user activity levels, and evaluate the model’s robustness when incorporating noise-based privacy enhancement. These analyses collectively highlight the significance of FedMR’s adaptive multimodal fusion and illustrate the interplay between modality utilization, user diversity, and privacy considerations in FedRec.
Figure 2. Performance comparison of different feature fusion strategies in various FedRecs combined with FedMR on KU.
5.6.1. Ablation Study on Fusion Strategy
Fig. 2 shows the HR@50 and NDCG@50 results for five FedRec methods with different fusion strategies on the KU dataset. The results highlight the critical role of fusion strategies in performance. Among them, the Mix strategy, which dynamically combines fusion methods based on user interaction history, achieves the best performance. In contrast, the Sum strategy performs the worst across all metrics. The Sum strategy, which simply adds features, fails to account for complex interactions among modalities, often leading to information loss and poor handling of long-tail distributions. In contrast, the MLP strategy better captures nonlinear interactions but is limited by its network depth and parameter count. The Gate strategy dynamically learns feature weights, optimizing the importance of different modalities. However, the Mix strategy combines all fusion methods, dynamically adjusting their contributions, enabling it to generate fine-grained personalized item representations and better capture complex relationships among features. This study underscores the importance of personalized item representations in FedMMRec. By leveraging the MFFM, FedMR allows each client to dynamically adjust fusion weights based on user interaction history, enabling fine-grained personalization and enhancing recommendation accuracy while maintaining privacy under the federated settings.
Table 3. Performance comparison of FedRecs with and without multimodal and ID features on KU.
Method Metric FedMR w/o V w/o C w/o D
@@ 5.6.1. Ablation Study on Fusion Strategy
FCF HR@50 0.5539 0.2843 (48.67% $\downarrow$) 0.4216 (23.89% $\downarrow$) 0.2353 (57.52% $\downarrow$)
NDCG@50 0.1597 0.0748 (53.16% $\downarrow$) 0.0965 (39.57% $\downarrow$) 0.0578 (63.81% $\downarrow$)
@@
PFedRec HR@50 0.2157 0.1990 (7.74% $\downarrow$) 0.1941 (10.01% $\downarrow$) 0.1794 (16.83% $\downarrow$)
NDCG@50 0.0616 0.0469 (23.86% $\downarrow$) 0.0537 (12.82% $\downarrow$) 0.0473 (23.21% $\downarrow$)
FedRAP HR@50 0.3824 0.2647 (30.78% $\downarrow$) 0.3235 (15.40% $\downarrow$) 0.1863 (51.29% $\downarrow$)
NDCG@50 0.1305 0.0762 (41.61% $\downarrow$) 0.0714 (45.29% $\downarrow$) 0.0490 (62.43% $\downarrow$)
@@ 5.6.2. Ablation Study on Modality Efficiency
We assess the contribution of visual, textual, and ID feature types by evaluating FedMR-enhanced model performance upon removing each source individually, referencing the updated results in Table 3. The findings reveal a clear interdependence among these modalities, confirming their complementary roles. ID features remain foundational for personalization, and their removal consistently causes the most significant performance degradation across baseline methods, often proving the most critical factor, particularly impacting models with strong personalization components like FedRAP. Visual features are also highly influential, with their absence substantially hindering performance, especially for visually-reliant architectures such as FCF. Textual features provide valuable semantic context, and removing them leads to notable accuracy decreases for various models, confirming their complementary contribution beyond visual and ID signals. Crucially, we can see the degree of reliance on each modality differs markedly across the underlying FedRec architectures. This differential dependency and demonstrated complementarity across modalities necessitate adaptive fusion mechanisms. FedMR’s MFFM addresses this requirement, ensuring robust performance by effectively utilizing all available information sources.
(a) User Distribution
(b) Modality Preferences
(c) Heterogeneity Distribution
(d) Performance Trend
Figure 3. Analysis of user characteristics, heterogeneity, and performance for FedRAP w/ FedMR on the KU dataset. (a) Distribution of users across interaction counts; (b) Variation in user preferences across modalities relative to the global average; (c) Distribution of user heterogeneity, which deviation from global preference, across interaction groups; (d) Recommendation performance trend on Recall@50 and NDCG@50 by user interaction group.
5.6.3. Ablation Study on Client Heterogeneity
We analyze client heterogeneity and its performance implications using the FedRAP enhanced by FedMR on the KU dataset, with results consolidated in Fig. 3. Fig. 3(a) shows that the dataset comprises users with varying activity levels. A key finding is the significant inter-user heterogeneity in modality preferences, as shown in Fig. 3(b). While preferences related to ID features exhibit relatively low variance, user preferences for text and especially vision features vary considerably more, highlighting the diverse ways users perceive and value different types of multimodal information, posing a challenge for standard fusion approaches. Furthermore, as we can see in Fig. 3(c), analyzing the overall heterogeneity distribution reveals that its variance tends to increase with user activity. More active users, therefore, not only deviate from the average preference but also exhibit greater diversity amongst themselves compared to less active users, suggesting that experienced users often develop more distinct and varied multimodal preference patterns, consistent with prior work (Liu et al., 2019; Guo et al., 2024).
This observed trend of increasing heterogeneity variance among more active users directly correlates with the recommendation performance shown in Fig. 3(d). Recall@50 and NDCG@50 tend to decrease for user groups with higher interaction counts. The heightened preference diversity within these active groups makes accurate Top-K prediction intrinsically more difficult. Capturing the varied, potentially niche multimodal preferences of highly heterogeneous users presents a significant challenge, contributing to the observed degradation in average performance metrics for these groups. This challenge underscores the limitations of non-adaptive fusion methods and highlights the core contribution of FedMR. Our framework introduces the MFFM, an innovative client-side mechanism specifically designed for dynamic, personalized adaptation of multimodal representations. By allowing each client to tailor the fusion process based on individual interaction history, MFFM directly addresses the crucial challenge of modality preference heterogeneity observed empirically, offering a vital step towards effective personalization in diverse federated multimodal environments.